[Ncep.nhc.nco_contacts] TSB Morning Rounds - Mon Jun 20, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Mon Jun 20 13:20:32 UTC 2016


Here is today's summary of NHC computer operations:
--- Mark

----------

TAFB:
This issue with the AWIPS EDEX restarting late last week was resolved.  
Chris Mello restarted these remotely as part of a development project. In  
the future, TSB will make sure to inform TAFB when the EDEX systems are  
restarted.

HSU:
Upper air data was not available in ntrans due to a disk crash on Mike  
Brennan's workstation. TSB was already in the process of migrating all  
operational processing from Mike's workstation, and this will continue to  
make sure critical processing is on the operationally supported compute  
server and its backup.

TSB is still investigating the spurious yellow vertical line on the ATCF  
when creating the TWO.

Dave resolved the problem with submitting a twitter message that occurred  
over the weekend. It was a minor syntax error that got introduced when we  
temporarily had to switch servers.

An unofficial "test" pre-TCM message and possibly a resend of the regular  
TCM went out yesterday. TSB is investigating.



Other items:

NHC RFCs:
RFC #179 - Adding a new script to download the experimental Spanish TCD  
(tropical cyclone discussion) since San Juan cannot transmit it via AWIPS.  
The script checks for active storms, then if there are any active storms,  
it compares the existing file with the one on the San Juan WFO website to  
see if it has been updated. If it has, the new file is send to the NHC web  
servers. To be implemented on June 20 at 1200Z.

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160620/a451f598/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list